USING THE EXTRAORDINARY INFRARED TRANSMISSION OF SURFACE PLASMON-MEDIATED METALLIC MICROARRAYS: INFRARED ABSORPTION ASSAY OF THE CATALYIC REACTION OF BENZYL ALCOHOL TO BENZYL BENZOATE ON NICKEL

\underline{KENNETH R. RODRIGUEZ}, DANIEL S. MELEASON, SHAUN M. WILLIAMS, SHAN-NON TEETERS-KENNEDY, KATHERINE CILWA, and JAMES V. COE, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

IR absorption has been recorded for the conversion of benzyl alcohol to benzyl benzoate on a nickel oxide surface by surface plasmon-mediated resonance. By activating the surface with OH groups, benzyl alcohol undergoes several intermediate steps that are recorded in time with benzyl benzoate forming the final product. Surface plasmon-mediated transmission resonance is a very sensitive technique that reveals dramatics changes at the metal-dielectric interface for the chemistry involved. Molecular absorption in the transmission spectra were analyzed to reveal those intermediate steps while obtaining an enhanced infrared absorption spectrum of benzyl benzoate.